Consider the two vectors, A = 18i + 39j B = 7.7i - 21.5j

What is the magnitude of the vector A - B? What is the direction of the vector A - B?
Answer #1

À = 18 i +39) B = 7.71 - 21.5 A - B = (18-7.1) + (39 +2 1.5 ] } - 10.39 + 60.5; magnitude of 1Ã-B1 = 1 10.3)2 + (60.5² = [61.